# A Survey of Vibe Coding with Large Language Models

### v1

@@ 9.4 Human Factors in Vibe Coding
In contrast to traditional programming where developers directly manipulate code logic, Vibe Coding emphasizes context engineering and human-AI collaboration [28, 623]. We examine three key aspects of this transformation: the shift in developers’ mental models, the evolution of required skill sets, and the reconfiguration of team collaboration and responsibility [624].
@@ 9.4.1 Mental Model Shift: From Code Logic to Context Engineering
Software engineering has historically been conceptualized as the process of translating requirements into algorithms and data structures expressed through code [625]. In Vibe Coding, however, developers increasingly act as context engineers, where the central task involves curating and structuring prompts, providing background information, and defining system constraints that guide AI-generated output [31, 626, 28]. Recent studies describe this shift as the emergence of prompt engineering [627, 628]. Rather than solving low-level implementation details directly, developers are expected to design high-level problem frames, construct testing scaffolds, and supply contextual cues, followed by iterative refinement of AI outputs [629].
Empirical studies on GitHub Copilot and similar tools indicate that developers often adopt a “specify–verify–revise” workflow, where prompting substitutes for manual implementation, and verification becomes the central locus of human activity [630, 631, 632]. This transformation aligns software engineering more closely with socio-technical practices such as product design and requirements engineering, where the clarity of intent expression is as important as technical implementation skills [633].
@@ 9.4.2 Evolving Developer Skill Sets
  • Prompting and context design. Developers must acquire the ability to formulate prompts that encode both intent and constraints [31, 627, 28]. This includes modular prompt construction, retrieval-augmented prompting, and the maintenance of prompt libraries as formal engineering artifacts [628].
  • Task decomposition. Research demonstrates that LLM-based agents perform more effectively when complex problems are decomposed into smaller, explicitly defined subtasks [634, 629]. This approach resonates with TDD but shifts the emphasis from manual implementation to steering AI behavior [530].
  • Quality supervision and verification. Studies highlight persistent issues in AI-generated code, including subtle bugs, security vulnerabilities, and hallucinated APIs [584, 635, 585, 636]. Developers must develop skills in automated testing, static analysis, and formal verification techniques as mechanisms of oversight [637, 632].
  • Agent governance and security. In addition to technical supervision, developers are increasingly responsible for managing access control, execution privileges, and provenance tracking of AI-generated outputs [624]. Literature on AI safety emphasizes the importance of least-privilege principles and verifiable provenance [638].
These emerging skill demands reposition developers from code producers to supervisors and orchestrators of AI-driven workflows, aligning with calls for human-in-the-loop governance in software engineering [639, 640].
@@ 9.4.3 Team Collaboration and AI Integration
The growing autonomy of coding agents introduces new challenges to team dynamics and responsibility allocation [640, 641]. Several empirical studies suggest that AI tools are increasingly capable of assuming roles traditionally associated with junior developers, such as generating boilerplate code, producing documentation, or proposing test cases [642, 643, 633]. In this sense, AI systems can be regarded as quasi-team members that reshape patterns of collaboration.
However, this reconfiguration also raises concerns regarding accountability and trust [624]. While developers benefit from efficiency gains, they also report risks of over-reliance and automation bias [628, 35, 641]. Effective trust calibration, characterized by neither blind acceptance nor excessive skepticism, is therefore critical [644, 624]. Beyond trust, questions of responsibility remain pressing: if an AI-generated code fragment introduces vulnerabilities, responsibility could lie with the developer, the reviewer, or the AI provider, underscoring the need for clearer accountability frameworks [638].
Collaborative workflows also require adaptation [641]. Existing review processes designed for human-written code have been shown to be inadequate for the scale and pace of AI-generated code [638]. To address this limitation, researchers propose new mechanisms including automated verification pipelines, provenance watermarks, and the integration of formal verification methods [632]. These innovations aim to balance the benefits of automation with the cognitive limits of human oversight [632, 640].
@@ 9.4.4 Broader Implications
The implications of Vibe Coding extend beyond team-level practices to education and organizational structures [641]. Scholars have argued that computing curricula should integrate training in prompting, AI governance, and human-AI collaboration to ensure that graduates acquire the competencies necessary to manage AI-augmented workflows [645, 28]. Without such curricular reform, future developers may lack the skills required for effective oversight [624].
At the organizational level, the adoption of AI-augmented development may lead to a reconfiguration of professional hierarchies [633]. Activities such as product design, requirements engineering, and test creation may gain in importance, while low-level code implementation becomes increasingly automated [643]. This development echoes earlier predictions that programmers of the future may function more as designers and supervisors than as traditional coders [646, 243, 640].
@@ 10 Conclusion
This paper advances Vibe Coding from scattered practice to a principled, research-grounded discipline. We formalize Vibe Coding as a triadic system that couples human intent and quality control ("what/why"), project context ("where"), and coding agents’ decision policies ("how"), providing a first constrained MDP formulation that specifies roles, interfaces, and optimization targets for agentic software development. Building on this foundation, we consolidate the field into five development models—Unconstrained Automation, Iterative Conversational Collaboration, Planning-Driven, Test-Driven, and Context-Enhanced—that practitioners can compose to meet distinct risk, speed, and governance requirements, and articulate the corresponding contributions as theoretical grounding, actionable guidance, and a forward-looking research agenda. More broadly, our triadic formulation positions Vibe Coding as an instance of human-cyber-physical systems [647, 648], reflecting the convergence of human intelligence, autonomous computation, and physical software artifacts in modern development. This paradigm resonates with the human-machine-thing universe proposed in network big data research [649] and human-in-the-loop computing that emphasizes organic interaction among humans, machines, and data [650], further aligning with the emerging vision of human-machine intelligence integration as a fundamental transformation in research and development methodologies [651].

### v2

@@ 9.4 Human Factors in Vibe Coding
In contrast to traditional programming where developers directly manipulate code logic, Vibe Coding emphasizes context engineering and human-AI collaboration [mei2025survey, schneider2024exploring]. We examine three key aspects of this transformation: the shift in developers’ mental models, the evolution of required skill sets, and the reconfiguration of team collaboration and responsibility [sabouri2025trust].
@@ 9.4.1 Mental Model Shift: From Code Logic to Context Engineering
Software engineering has historically been conceptualized as the process of translating requirements into algorithms and data structures expressed through code [brooks1995mythical]. In Vibe Coding, however, developers increasingly act as context engineers, where the central task involves curating and structuring prompts, providing background information, and defining system constraints that guide AI-generated output [white2023prompt, reynolds2021prompt, mei2025survey]. Recent studies describe this shift as the emergence of prompt engineering [liu2023pre, zamfirescu2023johnny]. Rather than solving low-level implementation details directly, developers are expected to design high-level problem frames, construct testing scaffolds, and supply contextual cues, followed by iterative refinement of AI outputs [shen2023taskbench].
Empirical studies on GitHub Copilot and similar tools indicate that developers often adopt a “specify–verify–revise” workflow, where prompting substitutes for manual implementation, and verification becomes the central locus of human activity [guglielmi2025copilot, hawlitschek2023empirical, flynn2025using]. This transformation aligns software engineering more closely with socio-technical practices such as product design and requirements engineering, where the clarity of intent expression is as important as technical implementation skills [xiao2025ai].
@@ 9.4.2 Evolving Developer Skill Sets
  • Prompting and context design. Developers must acquire the ability to formulate prompts that encode both intent and constraints [white2023prompt, liu2023pre, mei2025survey]. This includes modular prompt construction, retrieval-augmented prompting, and the maintenance of prompt libraries as formal engineering artifacts [zamfirescu2023johnny].
  • Task decomposition. Research demonstrates that LLM-based agents perform more effectively when complex problems are decomposed into smaller, explicitly defined subtasks [ma2023ai, shen2023taskbench]. This approach resonates with TDD but shifts the emphasis from manual implementation to steering AI behavior [cui2025tests].
  • Quality supervision and verification. Studies highlight persistent issues in AI-generated code, including subtle bugs, security vulnerabilities, and hallucinated APIs [pearce2025asleep, zhong2025developer, fu2023security, agarwal2024codemirage]. Developers must develop skills in automated testing, static analysis, and formal verification techniques as mechanisms of oversight [councilman2025towards, flynn2025using].
  • Agent governance and security. In addition to technical supervision, developers are increasingly responsible for managing access control, execution privileges, and provenance tracking of AI-generated outputs [sabouri2025trust]. Literature on AI safety emphasizes the importance of least-privilege principles and verifiable provenance [ambati2023security].
These emerging skill demands reposition developers from code producers to supervisors and orchestrators of AI-driven workflows, aligning with calls for human-in-the-loop governance in software engineering [molison2025llm, takerngsaksiri2025human].
@@ 9.4.3 Team Collaboration and AI Integration
The growing autonomy of coding agents introduces new challenges to team dynamics and responsibility allocation [takerngsaksiri2025human, mohamed2025impact]. Several empirical studies suggest that AI tools are increasingly capable of assuming roles traditionally associated with junior developers, such as generating boilerplate code, producing documentation, or proposing test cases [song2024impact, vasiliniuc2023case, xiao2025ai]. In this sense, AI systems can be regarded as quasi-team members that reshape patterns of collaboration.
However, this reconfiguration also raises concerns regarding accountability and trust [sabouri2025trust]. While developers benefit from efficiency gains, they also report risks of over-reliance and automation bias [zamfirescu2023johnny, chen2021evaluating, mohamed2025impact]. Effective trust calibration, characterized by neither blind acceptance nor excessive skepticism, is therefore critical [sahai2025amazon, sabouri2025trust]. Beyond trust, questions of responsibility remain pressing: if an AI-generated code fragment introduces vulnerabilities, responsibility could lie with the developer, the reviewer, or the AI provider, underscoring the need for clearer accountability frameworks [ambati2023security].
Collaborative workflows also require adaptation [mohamed2025impact]. Existing review processes designed for human-written code have been shown to be inadequate for the scale and pace of AI-generated code [ambati2023security]. To address this limitation, researchers propose new mechanisms including automated verification pipelines, provenance watermarks, and the integration of formal verification methods [flynn2025using]. These innovations aim to balance the benefits of automation with the cognitive limits of human oversight [flynn2025using, takerngsaksiri2025human].
@@ 9.4.4 Broader Implications
The implications of Vibe Coding extend beyond team-level practices to education and organizational structures [mohamed2025impact]. Scholars have argued that computing curricula should integrate training in prompting, AI governance, and human-AI collaboration to ensure that graduates acquire the competencies necessary to manage AI-augmented workflows [beale2025computer, mei2025survey]. Without such curricular reform, future developers may lack the skills required for effective oversight [sabouri2025trust].
At the organizational level, the adoption of AI-augmented development may lead to a reconfiguration of professional hierarchies [xiao2025ai]. Activities such as product design, requirements engineering, and test creation may gain in importance, while low-level code implementation becomes increasingly automated [vasiliniuc2023case]. This development echoes earlier predictions that programmers of the future may function more as designers and supervisors than as traditional coders [shneiderman2022human, dong2025survey, takerngsaksiri2025human].
@@ 10 Conclusion
This paper advances Vibe Coding from scattered practice to a principled, research-grounded discipline. We formalize Vibe Coding as a triadic system that couples human intent and quality control ("what/why"), project context ("where"), and coding agents’ decision policies ("how"), providing a first constrained MDP formulation that specifies roles, interfaces, and optimization targets for agentic software development. Building on this foundation, we consolidate the field into five development models—Unconstrained Automation, Iterative Conversational Collaboration, Planning-Driven, Test-Driven, and Context-Enhanced—that practitioners can compose to meet distinct risk, speed, and governance requirements, and articulate the corresponding contributions as theoretical grounding, actionable guidance, and a forward-looking research agenda. More broadly, our triadic formulation positions Vibe Coding as an instance of human-cyber-physical systems [pan2024integration, liu2020human], reflecting the convergence of human intelligence, autonomous computation, and physical software artifacts in modern development. This paradigm resonates with the human-machine-thing universe proposed in network big data research [wang2014network] and human-in-the-loop computing that emphasizes organic interaction among humans, machines, and data [chegn2022thinking], further aligning with the emerging vision of human-machine intelligence integration as a fundamental transformation in research and development methodologies [li2024ai4r].